# Logarithmic Sobolev inequality in manifolds with nonnegative curvature via the ABP method

### v2

@@ 3. The case of submanifolds
In this section, we assume that $(M^{n+m},g)$, with $n\geq 2$ and $m\geq 1$, is a complete noncompact manifold with nonnegative sectional curvature, and we fix a base point $o\in M$. Let $\Sigma\subset M$ be a compact submanifold of dimension $n$, possibly with boundary. Let $p=2$ and let $f$ be a smooth positive function defined in $\Sigma\setminus\partial\Sigma$ that vanishes on $\partial\Sigma$. Let $\bar{D}$ denote the the Levi-Civita connection of $M$, and let $D^{\Sigma}$ denote the induced connection on $\Sigma$. The second fundamental form $B$ of $\Sigma$ is defined by
@@
It suffices to prove Theorem 1.3 for the case where $\Sigma$ is connected. Let $\alpha>0$ be a given parameter and $r_{o}=d(o,\Sigma)$. By scaling, we may assume that
(5) $\int_{\Sigma}f^{2}\log f^{2}\,d\mathrm{vol}-\frac{1}{\alpha}\int_{\Sigma}|D^{\Sigma}f|^{2}\,d\mathrm{vol}-\frac{1}{4\alpha}\int_{\Sigma}f^{2}|H|^{2}\,d\mathrm{vol}=0.$
By the connectedness of $\Sigma$, there exists a solution of the following Neumann boundary problem
(6) $\left\{\begin{aligned} &\mathrm{div}(f^{2}D^{\Sigma}u)=f^{2}\log f^{2}-\frac{1}{\alpha}|D^{\Sigma}f|^{2}-\frac{1}{4\alpha}f^{2}|H|^{2},&\text{ in }\Sigma\setminus\partial\Sigma,\\ &\langle D^{\Sigma}u,\nu\rangle=\frac{r}{\alpha},&\text{ on }\partial\Sigma,\end{aligned}\right$.
where $\nu$ is the outward unit normal vector field of $\partial\Sigma$ with respect to $\Sigma$ and $r>1+\alpha r_{o}$ is an arbitrarily constant. By standard elliptic regularity theory, we know that $u\in C^{2,\gamma}$ for each $0<\gamma<1$.
As in [5], we set $U:=\{x\in\Sigma\setminus\partial\Sigma:|Du(x)|<\frac{r}{\alpha}\}$ and
$E:=\{(x,y):x\in U,y\in T^{\perp}_{x}\Sigma,|D^{\Sigma}u(x)|^{2}+|y|^{2}<\frac{r^{2}}{\alpha^{2}}\}.$
@@
$\Sigma_{r}:=\{\xi\in M:d(x,\xi)<\frac{r^{2}}{\alpha}\text{ for all }x\in\Sigma\}$
@@
$|\det D\Phi_{r}(x,y)|\leq r^{n+m}\Big(\frac{1}{r}+\frac{\Delta_{\Sigma}u(x)-\langle H(x),y\rangle}{n}\Big)^{n}$
@@
$e^{-\alpha\frac{|d(x,\Phi_{r}(x,y))|^{2}}{r^{2}}}|\det D\Phi_{r}(x,y)|\leq r^{n+m}e^{\frac{n}{r}-n}f(x)^{2}$
@@
$\mathrm{div}(f^{2}D^{\Sigma}u)$ $=f^{2}\Delta_{\Sigma}u+2\langle fD^{\Sigma}f,D^{\Sigma}u\rangle$
$=f^{2}\log f^{2}-\frac{1}{\alpha}|D^{\Sigma}f|^{2}-\frac{1}{4\alpha}f^{2}|H|^{2}.$
@@
$=\log f(x)^{2}-\frac{|D^{\Sigma}f(x)|^{2}}{\alpha f(x)^{2}}-\frac{1}{4\alpha}|H(x)|^{2}-2\langle\frac{D^{\Sigma}f(x)}{f},D^{\Sigma}u(x)\rangle-\langle H(x),y\rangle$
$\leq\log f(x)^{2}+\alpha|D^{\Sigma}u(x)|^{2}-\frac{1}{4\alpha}|H(x)|^{2}-\langle H(x),y\rangle$
$=\log f(x)^{2}+\alpha(|D^{\Sigma}u(x)|^{2}+y^{2})-\Big|\sqrt{\alpha}y+\frac{1}{2\sqrt{\alpha}}H(x)\Big|^{2},$
@@
$e^{-\alpha\frac{|d(x,\Phi_{r}(x,y))|^{2}}{r^{2}}}|\det D\Phi_{r}(x,y)|\leq r^{n+m}e^{\frac{n}{r}-n}e^{-\big|\sqrt{\alpha}y+\frac{1}{2\sqrt{\alpha}}H\big|^{2}}f(x)^{2}.$
@@
$\leq r^{n+m}e^{\frac{n}{r}-n}\int_{\Sigma}f(x)^{2}\Big(\int_{T_{x}^{\perp}\Sigma}e^{-\big|\sqrt{\alpha}y+\frac{1}{2\sqrt{\alpha}}H(x)\big|^{2}}\,dy\Big)\,d\mathrm{vol}(x)$
$=r^{n+m}e^{\frac{n}{r}-n}\Big(\frac{\pi}{\alpha}\Big)^{\frac{m}{2}}\int_{\Sigma}f(x)^{2}\,d\mathrm{vol}(x).$
@@
$\int_{0}^{\frac{r^{2}}{\alpha}-r_{o}}e^{-\alpha\frac{(r_{o}+\rho)^{2}}{r^{2}}}(n+m)\theta\omega_{n+m}\rho^{n+m-1}\,d\rho\leq r^{n+m}e^{\frac{n}{r}-n}\Big(\frac{\pi}{\alpha}\Big)^{\frac{m}{2}}\int_{\Sigma}f(x)^{2}\,d\mathrm{vol}(x).$
Dividing the above formula by $r^{n+m}$ and let $r\to\infty$, we have
$\theta e^{n}\Big(\frac{\pi}{\alpha}\Big)^{\frac{n}{2}}\leq\int_{\Sigma}f(x)^{2}\,d\mathrm{vol}(x).$
Consequently,
$n+\frac{n}{2}\log\frac{\pi}{\alpha}+\log\theta\leq\log\Big(\int_{\Sigma}f^{2}\,d\mathrm{vol}\Big).$
Combining this inequality with (5), one can get
@@
Let $\int_{\Sigma}f^{2}\,d\text{vol}=1$, we have
@@
Minimizing the right-hand side with respect to $\alpha>0$ yields
@@
$\int_{\Sigma}f^{2}\log f^{2}\leq\frac{n}{2}\log\Big(\frac{2}{\pi en}\theta^{-\frac{2}{n}}\int_{\Sigma}|Df|^{2}+\frac{1}{4}f^{2}|H|^{2}\,d\mathrm{vol}\Big).$
@@ Acknowledgements
We would like to thank Professor Frank Morgan for his useful correspondence.

### v3

@@ 3. The case of submanifolds
In this section, we assume that $(M^{n+m},g)$, with $n\geq 2$ and $m\geq 1$, is a complete noncompact manifold with nonnegative sectional curvature, and we fix a base point $o\in M$. Let $\Sigma\subset M$ be a compact submanifold of dimension $n$, possibly with boundary. Let $p=2$ and let $f$ be a nonnegative smooth function defined on $\Sigma$ that vanishes on $\partial\Sigma$. Let $\bar{D}$ denote the the Levi-Civita connection of $M$, and let $D^{\Sigma}$ denote the induced connection on $\Sigma$. The second fundamental form $B$ of $\Sigma$ is defined by
@@
It suffices to prove Theorem 1.3 for the case where $\Sigma$ is connected. Let $\alpha>0$ be a given parameter, $r_{o}=\max\{d(o,x):x\in\Sigma\}$ and $r>1+\alpha r_{0}$ is an arbitrary constant. By scaling, we can assume that
(5) $\int_{\Sigma}h^{2}\log h^{2}\,d\mathrm{vol}-\frac{1}{\alpha}\int_{\Sigma}|D^{\Sigma}h|^{2}+\frac{1}{4}h^{2}|H|^{2}\,d\mathrm{vol}=\frac{r}{\alpha}\int_{\partial\Sigma}h^{2}\,da,$
where $h=c(f+1/r)$ for some $c>0$. By the connectedness of $\Sigma$, there exists a solution of the following Neumann boundary problem
(6) $\left\{\begin{aligned} &\mathrm{div}(h^{2}D^{\Sigma}u)=h^{2}\log h^{2}-\frac{1}{\alpha}|D^{\Sigma}h|^{2}-\frac{1}{4\alpha}h^{2}|H|^{2},&\text{ on }\Sigma,\\ &\langle D^{\Sigma}u,\nu\rangle=\frac{r}{\alpha},&\text{ on }\partial\Sigma,\end{aligned}\right$.
where $\nu$ is the outward unit normal vector field of $\partial\Sigma$ with respect to $\Sigma$. By standard elliptic regularity theory, we know that $u\in C^{2,\gamma}$ for each $0<\gamma<1$.
As in [5], we set $U:=\{x\in\Sigma\setminus\partial\Sigma:|Du(x)|<r/\alpha\}$ and
$E:=\{(x,y):x\in U,y\in T^{\perp}_{x}\Sigma,|D^{\Sigma}u(x)|^{2}+|y|^{2}<r^{2}/\alpha^{2}\}.$
@@
$\Sigma_{r}:=\{\xi\in M:d(x,\xi)<r^{2}/\alpha\text{ for all }x\in\Sigma\}$
@@
$|\det D\Phi_{r}(x,y)|\leq r^{n+m}\bigg(\frac{1}{r}+\frac{\Delta_{\Sigma}u(x)-\langle H(x),y\rangle}{n}\bigg)^{n}$
@@
$e^{-\alpha\frac{d(x,\Phi_{r}(x,y))^{2}}{r^{2}}}|\det D\Phi_{r}(x,y)|\leq r^{n+m}e^{\frac{n}{r}-n}e^{-\big|\sqrt{\alpha}y+\frac{1}{2\sqrt{\alpha}}H\big|^{2}}h(x)^{2}$
@@
$\mathrm{div}(h^{2}D^{\Sigma}u)$ $=h^{2}\Delta_{\Sigma}u+2\langle hD^{\Sigma}h,D^{\Sigma}u\rangle$
$=h^{2}\log h^{2}-\frac{1}{\alpha}|D^{\Sigma}h|^{2}-\frac{1}{4\alpha}h^{2}|H|^{2}.$
@@
$=\log h(x)^{2}-\frac{|D^{\Sigma}h(x)|^{2}}{\alpha h(x)^{2}}-\frac{1}{4\alpha}|H(x)|^{2}-2\frac{\langle D^{\Sigma}h(x),D^{\Sigma}u(x)\rangle}{h(x)}-\langle H(x),y\rangle$
$\leq\log h(x)^{2}+\alpha|D^{\Sigma}u(x)|^{2}-\frac{1}{4\alpha}|H(x)|^{2}-\langle H(x),y\rangle$
$=\log h(x)^{2}+\alpha(|D^{\Sigma}u(x)|^{2}+y^{2})-\Big|\sqrt{\alpha}y+\frac{1}{2\sqrt{\alpha}}H(x)\Big|^{2},$
@@
$e^{-\alpha\frac{d(x,\Phi_{r}(x,y))^{2}}{r^{2}}}|\det D\Phi_{r}(x,y)|\leq r^{n+m}e^{\frac{n}{r}-n}e^{-\big|\sqrt{\alpha}y+\frac{1}{2\sqrt{\alpha}}H\big|^{2}}h(x)^{2}.$
@@
$\leq r^{n+m}e^{\frac{n}{r}-n}\int_{\Sigma}h(x)^{2}\Big(\int_{T_{x}^{\perp}\Sigma}e^{-\big|\sqrt{\alpha}y+\frac{1}{2\sqrt{\alpha}}H(x)\big|^{2}}\,dy\Big)\,d\mathrm{vol}(x)$
$=r^{n+m}e^{\frac{n}{r}-n}\Big(\frac{\pi}{\alpha}\Big)^{\frac{m}{2}}\int_{\Sigma}h(x)^{2}\,d\mathrm{vol}(x).$
@@
$\int_{0}^{\frac{r^{2}}{\alpha}-r_{o}}e^{-\alpha\frac{(r_{o}+\rho)^{2}}{r^{2}}}(n+m)\theta\omega_{n+m}\rho^{n+m-1}\,d\rho\leq r^{n+m}e^{\frac{n}{r}-n}\Big(\frac{\pi}{\alpha}\Big)^{\frac{m}{2}}\int_{\Sigma}h(x)^{2}\,d\mathrm{vol}(x).$
Combining this inequality with (5) and letting $h=f+\frac{1}{r}$ as $r\to\infty$, one can get
@@
Let $\int_{\Sigma}f^{2}\,d\text{vol}=1$, we find
@@
Minimizing the right hand side with respect to $\alpha>0$ yields
@@
$\int_{\Sigma}f^{2}\log f^{2}\leq\frac{n}{2}\log\bigg(\frac{2}{\pi en}\theta^{-\frac{2}{n}}\int_{\Sigma}|Df|^{2}+\frac{1}{4}f^{2}|H|^{2}\,d\mathrm{vol}\bigg).$
@@ Acknowledgements
We are grateful to Professor Frank Morgan for his useful correspondence, and to Professor Yimin Chen for his helpful discussions.